# Elasto-frictional reduced model of a cyclically sheared container filled with particles

### v1

@@ 3 Results
3.1 Evolution of the shear stress
@@
Once the sample is set under pressure at a given packing fraction, a quasi-harmonic simple shear strain shown in Fig. 1a is imposed to it as described in Sec. 2. The data on Fig. 1b present the evolution of the shear stress reaction between the walls of the container and the granular sample to accommodate such a motion, for two examples, at packing fraction $\phi=0.64$ and $\phi=0.70$. At such confinements, the average strain of the contacts (Roux et al., 2020) is $\langle h/a\rangle\approx 0.02$ and $\langle h/a\rangle\approx 0.04$, respectively, see the Fig. A1 in the Appendix Sec. A1. In both cases, a nonlinear response is clearly evidenced in the stress-strain relation $\tau(\gamma)$, see Figs. 1d and 1e. Indeed, we observe an hysteretic behavior on Fig. 1d at $\phi=0.70$, with a noticeable dependence with the shear amplitude $\gamma_{m}$. For instance, following clockwise the evolution of the stress-strain curve $\tau(\gamma)$ with the largest amplitude, shown as the darkest blue curve on Fig. 1d, starting from the left-bottom cusp point at $\gamma(t)=-0.15$, we can see that after a nonlinear growth, $\tau$ reaches a saturating point where it starts to increase linearly with respect to $\gamma$. Then $\tau$ reaches the right-top cusp point at $\gamma(t)=0.15$ before pursuing the same track backward. The behavior at $(\phi,\gamma_{m})=(0.70,0.15)$ has to be contrasted to the one at $(\phi,\gamma_{m})=(0.64,0.15)$ presented on Fig. 1e, see the darkest red curve: the stress-strain curve follows a similar trend, except that the linear growth before reaching cusp points becomes even more nonlinear, with an additional stiffening inflection. Such an inflection is less obvious in the case $(\phi,\gamma_{m})=(0.64,0.10)$ seen on Fig. 1e, for which the hysteresis curves appear more similar to the one at $(\phi,\gamma_{m})=(0.70,0.15)$. Also in both cases seen on Figs. 1d and 1e, the hysteresis tend to shrink and transform into a linear elastic response at the smallest amplitudes.
@@
All these observations thus highlight a symmetry of the roles played by the packing fraction, the confinement pressure and the shear strain amplitude on the mechanical response of the granular sample: large packing fractions, large confinement pressures and small strain amplitudes tend to linearize the stress-strain relation. This feature will be further discussed and exploited in more details in the following sections.
3.2 Evolution of the pressure
@@
The evolution of the pressure during cyclic shear is plotted on Fig. 1c. It shows that the pressure is fluctuating even though the cyclic shear transformation is isochoric. This phenomenon has the same origin as the so-called Reynolds dilatancy, an usual fluctuation of the volume fraction $\phi$ when the shear transformation is performed at constant pressure, as for instance in (Xing et al., 2021). In the later case, the granular medium has the tendency to expand when submitted to shear strain. In the case where the volume is constrained, the volume fraction is prescribed, the dilatancy is thus frustrated, which results in a fluctuation of the pressure. This variation, called the Reynolds pressure in (Ren et al., 2013), can be described as
@@
where $P$, $P_{0}$ and $\delta P$ correspond to the time average, the static value and the fluctuation of the pressure, respectively. Because of the quadratic dependency in shear deformation, the static pressure $P_{0}$ (i.e. the pressure at zero strain, $\gamma_{m}=0$) and the mean pressure $P=P_{0}+\langle R\gamma^{2}\rangle$ are not equal; however, they remain close for small strain amplitudes, $R\gamma_{m}^{2}\ll P$. The Reynolds coefficient $R$ was first introduced by Ren et al in (Ren et al., 2013), where it was demonstrated to be related to the volume fraction, $R=A(\phi_{c}-\phi)^{n}$ with $n\simeq-3.3$ for cylindrical particles in 2D and $A$ having the dimension of an elasticity. It is worth pointing out that the behavior of $R$ was examined below jamming in (Ren et al., 2013), $\phi<\phi_{c}$, while our study is conducted in a 3D packing of spherical grains above jamming, $\phi>\phi_{c}$, which is likely to provide a different scaling of $R$ with respect to the packing fraction. In our case, we find that $R\propto(\phi-\phi_{c})^{n}$ with $n\simeq 0.56$, see Fig. A2a in the Appendix Sec. A2. In an alternative manner, a dependency of the pressure $P$ with the shear strain $\gamma$ was also captured by the constitutive model derived by Luding et al in (Krijgsman and Luding, 2013), which demonstrates that the pressure fluctuations in a 2D granular material under cyclic pure shear result from a structural anisotropy, coupling volumic and deviatoric component in the stress tensor.
@@
In Fig. 2, we check the relevance of Eq. 3 for the two examples shown in Fig. 1. For this purpose, we introduce the peak amplitude of the pressure fluctuations, $\Delta P=R\gamma_{m}^{2}/2$, stemming from the quadratic dependency $\gamma^{2}(t)=\gamma_{m}^{2}\sin^{2}{(\omega_{0}t)}=(\gamma_{m}^{2}/2)[1-\cos{(2\omega_{0}t)}]$. At first, the Figs. 2a and 2b reveal that the spectral component of the normalized pressure fluctuation $\delta P/\Delta P$ is mostly located at twice the driving frequency $f_{0}$ prescribed by the quasi-harmonic shear strain $\gamma/\gamma_{m}$. Such a frequency conversion is typical of a quadratic dependency, consistently with the reliable quadratic fit shown in Figs. 2c and 2d. The estimated values of $R$ from these fits, $R=4.49$ GPa for $(\phi,P)=(0.64,257\mathrm{~{}MPa})$ and $R=7.54$ GPa for $(\phi,P)=(0.70,929\mathrm{~{}MPa})$, indicate that $R$ is a growing function of $P$ and $\phi$. In more details, the Appendix Sec. A2 reveals a power law of the pressure, $R\propto P^{m}$ with $m\simeq 0.37$, see Fig. A2b, consistently with the aforementioned power law of the packing fraction, $R\propto(\phi-\phi_{c})^{n}$ with $n\simeq 0.56$, see Fig. A2a.
Finally, we observe on Fig. 1c that the lower the packing fraction or the pressure, the greater $\delta P/P$. This observation is compatible with the evolution of the nonlinear behavior evidenced in Figs. 1d and 1e: the pressure being a state variable, its fluctuation with the shear deformation will likely make the mechanical response nonlinear. Indeed, the elasticity of granular media relies on Hertz contact theory, see above in Sec. 2, which depends on the confining pressure, see below in Sec. 4.1: a significant fluctuation of the pressure implies a nonlinearity of the elasticity. At first in the following, we will thus restrict our description to cases at sufficiently small $\delta P/P$, i.e. at large packing fractions and pressures, in order to deal with linear elasticity; next, we will account for additional nonlinear contributions at smaller packing fractions and pressures.
@@ 3.3 Elasto-hysteretic behavior
Here, we aim at reducing the complexity of the description of numerous particles submitted to shear, by identifying a low order approximation depending on essential parameters only. To reveal this model, we first focus on the example with weak relative fluctuation of $P$, at $(\phi,\gamma_{m})=(0.70,0.15)$ from Fig. 1d. The stress-strain relation of this example is reproduced in Fig. 3. At first glance, one sees that such a mechanical response is likely to be decomposed with a hysteretic response (a path-dependent saturating contribution) superimposed with a proportional response (a straight line with constant slope). In the following, we propose to describe the former with the Dahl model (Dahl, 1968, 1977), while the latter is expressed as a linear elastic term. Such an elasto-hysteretic decomposition was also used to describe the response of fibrous materials submitted to shear deformation (Piollet et al., 2016). The Dahl model is classically used in control engineering as an empirical model due to its implementation facility; in particular, it was shown to constitute an alternative formulation of the Mindlin frictional interaction between a sphere and a plane (Peyret et al., 2017). The shear stress is thus expressed as
@@
where $G_{D}$, the first parameter of Dahl’s model, is homogeneous to a shear modulus and corresponds to the slope of $\tau_{D}(\gamma)$ when $\tau_{D}=0$. The second parameter, $\mu_{D}$, corresponds to a Coulomb friction coefficient so that $\mu_{D}P$ is the asymptotic value of $\tau_{D}$ at large shear strain. The third parameter, the exponent $n_{D}$, controls the shape of the curve, e.g. $n_{D}<1$ exhibits brittle behaviors, while $n_{D}>1$ renders ductile responses. In the following of the study, we choose $n_{D}=1$ as a first attempt, in line with the literature (Piollet et al., 2016; Helmick and Messner, 2009; Al Majid and Dufour, 2004), to model elasto-plastic behaviors in between ductile and fragile.
The parameter identification consists in finding the optimal parameters triplet $(\mu_{D}^{\mathrm{opt}},G_{D}^{\mathrm{opt}},G_{E}^{\mathrm{opt}})$ that minimizes the difference between the reduced model defined in Eqs. 4 and 5, and the steady mechanical response simulated with DEM. More details on the process of identification is given in the Appendix Sec. A3. To demonstrate the reliability of such identification, we show the root mean square error (RMSE) between $\tau(\mu_{D},G_{D},G_{E})$ and $\tau_{\mathrm{DEM}}$ across three cross-sections of the parameters space, see Figs. 3a, 3b, and 3c. One can see that an unique and well-defined global minimum exists in a wide region of parameters. The qualitative agreement of the fit is shown on Fig. 3d for the example $(\phi,\gamma_{m})=(0.70,0.15)$, and the quantitative agreement is illustrated in ore details in the Appendix Sec. A3 on Fig. A3.
Figure 3: (a,b,c) Maps of the RMSE between $\tau(\mu_{D},G_{D},G_{E})$ and $\tau_{\mathrm{DEM}}$ plotted in the three planes $(\mu_{D},G_{E})$, $(\mu_{D},G_{D})$ and $(G_{D},G_{E})$, for the example $(\phi,\gamma_{m})=(0.70,0.15)$ shown in Fig. 1d. The parameters are scaled by the optimal triplet $\mu_{D}^{\mathrm{opt}}=0.24$, $G_{D}^{\mathrm{opt}}=9.7$ GPa and $G_{E}^{\mathrm{opt}}=2.3$ GPa. (d) Stress-strain relation from DEM simulation superimposed to the fits. The purple dotted line represent the hysteretic Dahl contribution $\tau_{D}(\gamma)$, the solid blue line represents the linear contribution $G_{E}\gamma$, see Eq. 4, and the red dashed line is the sum of the two.
@@ 3.4 Energy dissipation
The hysteresis of the macroscopic stress-strain relation evidenced in Fig. 3 renders the mechanical response non-conservative because of a path-dependent mechanical loading/unloading. Likely being the fingerprint of the microscopic inter-particle friction inside the granular sample, such a mechanism thus introduces energy dissipation; we aim in this subsection at defining and quantifying the damping properties of our system. In the following, the dissipated energy per volume unit and per cycle is measured as
[…]
@@
As an example, the definitions Eqs. 6-8 applied to the simulated strain-stress relation shown in Fig. 3d indicate that dissipated energy per cycle is $E_{\mathrm{diss}}^{\mathrm{cycle}}\simeq 114$ J/cm3, whereas the maximum stored energy is $E_{\mathrm{stored}}^{\mathrm{max}}\simeq 45$ J/cm3, leading to a loss factor $\eta\simeq 0.40$. Such a level of damping is worthy of interest in engineering applications. For sake of comparison, an efficient viscoelastic material such as Deltane 350 from Paulstra manufacturer, a synthetic elastomer commonly used in the aeronautics field, ranges from lower value $\eta\simeq 0.16$ at $1$ Hz to a higher value $\eta\simeq 0.88$ at $100$ Hz, reaching $\eta\simeq 0.45$ at $10$ Hz (Rouleau, 2013, p.94). In contrast with the frequency-independent friction-based dissipation observed in our granular medium, the polymer material thus loses its efficiency at low frequency, in addition to being strongly temperature dependent (Rouleau, 2013, p.76).
@@ 3.5 Parametric study
We now look at generalizing our observations for various configurations in the parameter space. In particular, we seek at elucidating the possibility of an optimal regime for energy dissipation, in between low confinement pressure (where the particles freely flow on top of each other without rubbing) and high confinement pressure (where the particles firmly stick to form a sturdy elastic brick). Three families of parameters are distinguished, namely (i) the particle properties, (ii) the state variables defining the packing and (iii) the external parameters defining the driving excitation.
@@
The particle properties include their Young modulus $E_{p}$, Poisson ratio $\nu_{p}$, mass density $\rho_{p}$, radius $r_{p}$, number $N_{p}$ and the inter-particle friction coefficient $\mu_{p}$. For sake of simplicity, fictional materials are considered, for which those parameters are varied independently; however, they range within realistic values, as for the Young modulus $E_{p}$ reported in Tab. 1. Without loss of generality, we assume that the Poisson ratio has a negligible effect (it only contributes to a second order correction of the elasticity, see Eqs. 1 and 2) so that it is arbitrarily kept constant in this study, equal to its nominal value $\nu_{p}=0.27$. Also, the mechanical response should be independent of the mass of the particles at sufficiently low driving frequency (see below, in the quasi-static regime), the mass density considered in the following parametric study has a nominal value $\rho_{p}=7850$ kg/m3. In the same manner, it is likely that the size and the number of particles should not play a role at the macroscopic scale, in the long wavelength (low frequency, see below) continuum limit approximation. The insensitivity of the mass, the size and the number of the particles is probed at once successfully by simulating a sample made of particles with half the average radius (i.e. one eighth of the mass per particle) but same sample’s volume (i.e. eight times more particles) without noticeable difference compared to the nominal case; this case is shown in the Sec. 4.1. Finally, the effect of inter-particle friction coefficient $\mu_{p}$ is probed separately, over more than a decade, in Sec. 4.2.
@@
The state variables include the packing fraction $\phi$, the pressure $P$ and the coordination number $Z$, see Tab. 1. These three quantities are related: according to the litterature (O’Hern et al., 2003), we checked (result not shown, see (Faulconnier, 2023, p.111)) that $P\propto(\phi-\phi_{c})^{3/2}$ and $(Z-Z_{c})\propto(\phi-\phi_{c})^{1/2}$ where $\phi_{c}=0.595\pm 0.003~{}(0.5\%)$ and $Z_{c}=5.042\pm 0.085~{}(1.7\%)$ are values of the packing fraction and coordination number at the critical point where the jamming occurs. Thus, imposing the packing fraction $\phi$ sets the average level of pressure $P$ and the average contact number $Z$.
@@
Finally, the external parameters include the excitation angular frequency $\omega_{0}=2\pi f_{0}$ and shear amplitude $\gamma_{m}$, see Tab. 1. The former is kept low enough, to fulfill a quasi-static regime where the inertia of the grains does not produce any contribution in the force balance. The latter is kept constant, in the parametric study presented in Secs. 4.1 and 4.2, at a nominal value $\gamma_{m}=0.15$ that ensures observing noticeable hysteresis in the stress-strain relation, as in Figs. 1d and 1e for instance. The dataset is then completed by probing different amplitudes, in Secs. 4.3 and 4.4.
@@
$\gamma_{m}$ $0.005\sim 0.15$
For each configuration of the parameters space, the system is simulated using DEM, from which we estimate the stress-strain curve in shear as described in Sec. 2. Then, an optimal triplet of parameters $(\mu_{D},G_{D},G_{E})$ relying on the elasto-hysteretic reduced model are retrieved according to the protocol described in Sec. 3.3. These three parameters are plotted versus the state variable $P$ in Fig. 4a and 4b for a wide range of particle properties. For dimensional reasons, we normalize the pressure $P$ and the elasticity $G_{E,D}$ by the inter-particle contact stiffness (proportional to the Young’s modulus of the particles material) in the normal direction, $k_{n}\propto E_{p}$, as defined in Eq. 1. One can see on Fig. 4a that both $G_{E}$ and $G_{D}$ increase with $P$, over a decade, resulting in two clearly defined but distinct slopes which will be analyzed in-depth in the Sec. 4. In contrast, the effective Coulomb friction coefficient $\mu_{D}$ shown in Fig. 4b appears weakly dependent of the confinement pressure and of the particle properties, with an average value $\langle\mu_{D}\rangle\simeq 0.28$.
Finally, we look at the energetical performances of the granular packing, as defined in Eqs. 6-8; again, we normalize the energies per unit volume by $k_{n}$ (i.e. by the particle’s Young Modulus) for dimensional reasons. One sees on Fig. 4c that both the dissipated and the maximum stored energy per cycle increase as functions of the confinement pressure, consistently with the elastic features of the effective model, $G_{E}$ and $G_{D}$. Also in agreement with the effective Coulomb friction coefficient $\mu_{D}$, their ratio (i.e. the loss factor), shown in Fig. 4d, appears weakly dependent of the pressure and the particle properties, lying close to the value $\eta=0.40$ already observed in Fig. 3d.
Figure 4: (a): Dimensionless fitted parameters $G_{D}/k_{n}$ (open symbols) and $G_{E}/k_{n}$ (solid symbols) for various volume fraction $\phi$ and elasticity $E_{p}$ as functions of the dimensionless pressure $P/k_{n}$ at $\gamma_{m}=0.15$. (b): Dahl friction coefficient against dimensionless pressure for the same data set. (c): Dissipated energy per cycle (open symbols) and maximum stored energy (solid symbols) computed from hysteresis response using Eqs. 6 and 7 respectively, as functions of the dimensionless pressure. (d): Loss factor computed from 8, as function of the dimensionless pressure. Legend symbols are common for (a-d).

### v2

@@ 3 Results
3.1 Evolution of shear stress
@@
Once the sample is set under pressure at a given packing fraction, a quasi-harmonic simple shear strain shown in Fig. 1a is imposed on it as described in Sec. 2. The data in Fig. 1b present the evolution of the shear stress reaction between the walls of the container and the granular sample to accommodate such a motion for two examples at packing fraction $\phi=0.64$ and $\phi=0.70$. At such confinements, the average strain of the contacts (Roux et al., 2020) is $\langle h/a\rangle\approx 0.02$ and $\langle h/a\rangle\approx 0.04$, respectively, see Fig. A1 in the Appendix Sec. A1. In both cases, a nonlinear response is clearly evidenced in the stress-strain relation $\tau(\gamma)$, see Figs. 1d and 1e. Indeed, we observe a hysteretic behavior in Fig. 1d at $\phi=0.70$, with a noticeable dependence on the shear amplitude $\gamma_{m}$. For instance, following the evolution of the stress-strain curve $\tau(\gamma)$ with the largest amplitude, shown as the darkest blue curve in Fig. 1d, in the clockwise direction by starting from the left-bottom cusp point at $\gamma(t)=-0.15$, we can see that after a nonlinear growth, $\tau$ reaches a saturating point where it starts to increase linearly with respect to $\gamma$. Then $\tau$ reaches the right-top cusp point at $\gamma(t)=0.15$ before pursuing the same track backward. The behavior at $(\phi,\gamma_{m})=(0.70,0.15)$ must be contrasted to that at $(\phi,\gamma_{m})=(0.64,0.15)$ presented in Fig. 1e, see the darkest red curve: the stress-strain curve follows a similar trend, except that the linear growth before reaching cusp points becomes even more nonlinear, with an additional stiffening inflection. Such an inflection is less obvious in the case $(\phi,\gamma_{m})=(0.64,0.10)$ seen in Fig. 1e, for which the hysteresis curves appear more similar to that at $(\phi,\gamma_{m})=(0.70,0.15)$. Also in both cases seen in Figs. 1d and 1e, the hysteresis tends to shrink and transform into a linear elastic response at the smallest amplitudes.
@@
All these observations thus highlight a symmetry of the roles played by the packing fraction, the confinement pressure and the shear strain amplitude on the mechanical response of the granular sample: large packing fractions, large confinement pressures and small strain amplitudes tend to linearize the stress-strain relation. This feature will be discussed further and exploited in more detail in the following sections.
3.2 Evolution of pressure
@@
The evolution of the pressure during cyclic shear is plotted on Fig. 1c. It shows that the pressure is fluctuating even though the cyclic shear transformation is isochoric. This phenomenon has the same origin as the so-called Reynolds dilatancy, a usual fluctuation of the volume fraction $\phi$ when the shear transformation is performed at constant pressure, as for instance in (Xing et al., 2021). In the latter case, the granular medium tends to expand when subjected to shear strain. In the case where the volume is constrained, the volume fraction is prescribed and the dilatancy is therefore restrained, resulting in a fluctuation of the pressure. This variation, called Reynolds pressure in (Ren et al., 2013), can be described as
@@
where $P$, $P_{0}$ and $\delta P$ correspond to the time average, the static value and the fluctuation of the pressure, respectively. Because of the quadratic dependency in shear deformation, the static pressure $P_{0}$ (i.e. the pressure at zero strain, $\gamma_{m}=0$) and the mean pressure $P=P_{0}+\langle R\gamma^{2}\rangle$ are not equal; however, they remain close for small strain amplitudes, $R\gamma_{m}^{2}\ll P$. The Reynolds coefficient $R$ was first introduced by Ren et al in (Ren et al., 2013), when it was demonstrated to be related to the volume fraction, $R=A(\phi_{c}-\phi)^{n}$ with $n\simeq-3.3$ for disk packings in 2D and $A$ having the dimension of an elasticity. It is worth pointing out that the behavior of $R$ was examined below jamming in (Ren et al., 2013), $\phi<\phi_{c}$, while our study is conducted in a 3D packing of spherical grains above jamming, $\phi>\phi_{c}$, which is likely to provide a different scaling of $R$ with respect to the packing fraction. In our case, we find that $R\propto(\phi-\phi_{c})^{n}$ with $n\simeq 0.56$, see Fig. A2a in the Appendix Sec. A2. In an alternative manner, a dependency of the pressure $P$ with the shear strain $\gamma$ was also captured by the constitutive model derived by Luding et al in (Krijgsman and Luding, 2013), which demonstrates that the pressure fluctuations in a 2D granular material under cyclic pure shear result from structural anisotropy, coupling volumetric and deviatoric components in the stress tensor.
@@
In Fig. 2, we check the relevance of Eq. 3 for the two examples shown in Fig. 1. To this end, we introduce the peak amplitude of the pressure fluctuations, $\Delta P=R\gamma_{m}^{2}/2$, stemming from the quadratic dependency $\gamma^{2}(t)=\gamma_{m}^{2}\sin^{2}{(\omega_{0}t)}=(\gamma_{m}^{2}/2)[1-\cos{(2\omega_{0}t)}]$. First, Figs. 2a and 2b reveal that the spectral component of the normalized pressure fluctuation $\delta P/\Delta P$ is mostly located at twice the driving frequency $f_{0}$ prescribed by the quasi-harmonic shear strain $\gamma/\gamma_{m}$. Such a frequency conversion is typical of a quadratic dependency, consistent with the reliable quadratic fit shown in Figs. 2c and 2d. The estimated values of $R$ from these fits, $R=4.49$ GPa for $(\phi,P)=(0.64,257\mathrm{~MPa})$ and $R=7.54$ GPa for $(\phi,P)=(0.70,929\mathrm{~MPa})$, indicate that $R$ is a growing function of $P$ and $\phi$. In greater detail, Appendix Sec. A2 reveals a power law of the pressure, $R\propto P^{m}$ with $m\simeq 0.37$, see Fig. A2b, consistent with the aforementioned power law of the packing fraction, $R\propto(\phi-\phi_{c})^{n}$ with $n\simeq 0.56$, see Fig. A2a.
Finally, we observe in Fig. 1c that the lower the packing fraction or the pressure, the higher $\delta P/P$ is. This observation is compatible with the evolution of the nonlinear behavior evidenced in Figs. 1d and 1e: the pressure being a state variable, its fluctuation with the shear deformation will likely make the mechanical response nonlinear. Indeed, the elasticity of granular media relies on Hertz contact theory, see above in Sec. 2, which depends on the confining pressure, see below in Sec. 4.1: a significant fluctuation of the pressure implies a nonlinearity of the elasticity. In the following, we will thus restrict our description to cases at sufficiently small $\delta P/P$, i.e. at large packing fractions and pressures, to deal with linear elasticity; next, we will account for additional nonlinear contributions at smaller packing fractions and pressures.
@@ 3.3 Elasto-hysteretic behavior
Here, we aim at reducing the complexity of the description of numerous particles subjected to shear, by identifying a low order approximation depending on essential parameters only. To determine this model, we first focus on the example with a weak relative fluctuation of $P$, at $(\phi,\gamma_{m})=(0.70,0.15)$ from Fig. 1d. The stress-strain relation of this example is reproduced in Fig. 3. At first glance, it can be seen that such a mechanical response is likely to be decomposed with a hysteretic response (a path-dependent saturating contribution) superimposed with a proportional response (a straight line with constant slope). In the following, we describe the former with the Dahl model (Dahl, 1968, 1977), while the latter is expressed as a linear elastic term. Such an elasto-hysteretic decomposition was also used to describe the response of fibrous materials subjected to shear deformation (Piollet et al., 2016). The Dahl model is classically used in control engineering as an empirical model due to its facility of implementation; in particular, it was shown to constitute an alternative formulation of the Mindlin frictional interaction between a sphere and a plane (Peyret et al., 2017). The shear stress is thus expressed as
@@
where $G_{D}$, the first parameter of Dahl’s model, is homogeneous with a shear modulus and corresponds to the slope of $\tau_{D}(\gamma)$ when $\tau_{D}=0$. The second parameter, $\mu_{D}$, corresponds to a Coulomb friction coefficient so that $\mu_{D}P$ is the asymptotic value of $\tau_{D}$ at large shear strain. The third parameter, exponent $n_{D}$, controls the shape of the curve, e.g., $n_{D}<1$ exhibits brittle behaviors, while $n_{D}>1$ renders ductile responses. In the rest of the study, we choose $n_{D}=1$ as the first attempt, in line with the literature (Piollet et al., 2016; Helmick and Messner, 2009; Al Majid and Dufour, 2004), to model elasto-plastic behaviors in between ductile and fragile.
The parameter identification consists in finding the optimal parameter triplet $(\mu_{D}^{\mathrm{opt}},G_{D}^{\mathrm{opt}},G_{E}^{\mathrm{opt}})$ that minimizes the difference between the reduced model defined in Eqs. 4 and 5, and the steady mechanical response simulated with DEM. More details on the process of identification is given in the Appendix Sec. A3. To demonstrate the reliability of such identification, we show the root mean square error (RMSE) between $\tau(\mu_{D},G_{D},G_{E})$ and $\tau_{\mathrm{DEM}}$ across three cross-sections of the parameters space, see Figs. 3a, 3b, and 3c. It can be seen that a unique and well-defined global minimum exists in a wide region of parameters. The qualitative agreement of the fit is shown in Fig. 3d for the example $(\phi,\gamma_{m})=(0.70,0.15)$, and the quantitative agreement is illustrated in more detail in the Appendix Sec. A3 in Fig. A3.
Figure 3: (a,b,c) Maps of the RMSE between $\tau(\mu_{D},G_{D},G_{E})$ and $\tau_{\mathrm{DEM}}$ plotted in the three planes $(\mu_{D},G_{E})$, $(\mu_{D},G_{D})$ and $(G_{D},G_{E})$, for the example $(\phi,\gamma_{m})=(0.70,0.15)$ shown in Fig. 1d. The parameters are scaled by the optimal triplet $\mu_{D}^{\mathrm{opt}}=0.24$, $G_{D}^{\mathrm{opt}}=9.7$ GPa and $G_{E}^{\mathrm{opt}}=2.3$ GPa. (d) Stress-strain relation from DEM simulation superimposed on the fits. The purple dotted line represents the hysteretic Dahl contribution $\tau_{D}(\gamma)$, the solid blue line represents the linear contribution $G_{E}\gamma$, see Eq. 4, and the red dashed line is the sum of the two.
@@ 3.4 Energy dissipation
The hysteresis of the macroscopic stress-strain relation evidenced in Fig. 3 renders the mechanical response non-conservative because of a path-dependent mechanical loading/unloading. Likely the fingerprint of the microscopic inter-particle friction inside the granular sample, this mechanism therefore introduces energy dissipation. In this subsection we aim at defining and quantifying the damping properties of our system. In the following, the dissipated energy per volume unit and per cycle is measured as
$E_{\mathrm{diss}}^{\mathrm{cycle}}=\frac{1}{n}\oint_{n~\mathrm{cycles}}\tau d\gamma,$ (6)
@@
As an example, the definitions Eqs. 6-8 applied to the simulated strain-stress relation shown in Fig. 3d indicate that the dissipated energy per cycle is $E_{\mathrm{diss}}^{\mathrm{cycle}}\simeq 114$ J/cm3, whereas the maximum stored energy is $E_{\mathrm{stored}}^{\mathrm{max}}\simeq 45$ J/cm3, leading to a loss factor $\eta\simeq 0.40$. Such a level of damping is worthy of interest in engineering applications. For the sake of comparison, an efficient viscoelastic material such as Deltane 350 from Paulstra, a synthetic elastomer commonly used in the aeronautics field, ranges from a lower value $\eta\simeq 0.16$ at $1$ Hz to a higher value $\eta\simeq 0.88$ at $100$ Hz, reaching $\eta\simeq 0.45$ at $10$ Hz (Rouleau, 2013, p.94). In contrast with the frequency-independent friction-based dissipation observed in our granular medium, the polymer material thus loses its efficiency at low frequency, in addition to being strongly temperature dependent (Rouleau, 2013, p.76).
@@ 3.5 Parametric study
We now look at generalizing our observations for various configurations in the parameter space. In particular, we seek to elucidate the possibility of an optimal regime for energy dissipation, in between low confinement pressure (where the particles freely flow on top of each other without rubbing) and high confinement pressure (where the particles firmly stick to form a sturdy elastic brick). Three families of parameters are distinguished, namely (i) the particle properties, (ii) the state variables defining the packing, and (iii) the external parameters defining the driving excitation.
@@
The particle properties include their Young’s modulus $E_{p}$, Poisson’s ratio $\nu_{p}$, mass density $\rho_{p}$, radius $r_{p}$, number $N_{p}$ and the inter-particle friction coefficient $\mu_{p}$. For the sake of simplicity, fictional materials are considered, for which these parameters vary independently; however, they range within realistic values, like for the Young’s modulus $E_{p}$ reported in Tab. 1. Without loss of generality, we assume that the Poisson’s ratio has a negligible effect (it only contributes to a second order correction of the elasticity, see Eqs. 1 and 2) so that it is arbitrarily kept constant in this study, equal to its nominal value $\nu_{p}=0.27$. Also, the mechanical response should be independent of the mass of the particles at sufficiently low driving frequency (see below, in the quasi-static regime), the mass density considered in the following parametric study has a nominal value $\rho_{p}=7850$ kg/m3. In the same manner, it is likely that the size and the number of particles should not play a role at the macroscopic scale, in the long wavelength (low frequency, see below) continuum limit approximation. The insensitivity of the mass, the size and the number of the particles is probed at once successfully by simulating a sample made of particles with half the average radius (i.e. one eighth of the mass per particle) but the same sample’s volume (i.e. eight times more particles) without a noticeable difference compared to the nominal case. This case is shown in Sec. 4.1. Finally, the effect of inter-particle friction coefficient $\mu_{p}$ is probed separately, over more than a decade, in Sec. 4.2.
@@
The state variables include the packing fraction $\phi$, the pressure $P$ and the coordination number $Z$, see Tab. 1. These three quantities are related: according to the literature (O’Hern et al., 2003), we checked (result not shown, see (Faulconnier, 2023, p.111)) that $P\propto(\phi-\phi_{c})^{3/2}$ and $(Z-Z_{c})\propto(\phi-\phi_{c})^{1/2}$ where $\phi_{c}=0.595\pm 0.003~(0.5\%)$ and $Z_{c}=5.042\pm 0.085~(1.7\%)$ are the values of the packing fraction and coordination number at the critical point where the jamming occurs. Thus, imposing the packing fraction $\phi$ sets the average level of pressure $P$ and the average contact number $Z$.
@@
Finally, the external parameters include the excitation angular frequency $\omega_{0}=2\pi f_{0}$ and shear amplitude $\gamma_{m}$, see Tab. 1. The former is kept low enough to fulfill a quasi-static regime where the inertia of the grains does not make any contribution to the force balance. The latter is kept constant in the parametric study presented in Secs. 4.1 and 4.2 at a nominal value $\gamma_{m}=0.15$ that ensures observing noticeable hysteresis in the stress-strain relation, as in Figs. 1d and 1e for instance. The dataset is then completed by probing different amplitudes, in Secs. 4.3 and 4.4.
@@
$\gamma_{m}$ $0.005\sim 0.15$ -
For each configuration of the parameter space, the system is simulated using DEM, from which we estimate the stress-strain curve in shear as described in Sec. 2. Then, an optimal triplet of parameters $(\mu_{D},G_{D},G_{E})$ relying on the elasto-hysteretic reduced model are retrieved according to the protocol described in Sec. 3.3. These three parameters are plotted versus the state variable $P$ in Fig. 4a and 4b for a wide range of particle properties. For dimensional reasons, we normalize the pressure $P$ and the elasticity $G_{E,D}$ by the inter-particle contact stiffness (proportional to the Young’s modulus of the particles’ material) in the normal direction, $k_{n}\propto E_{p}$, as defined in Eq. 1. Fig. 4a shows that both $G_{E}$ and $G_{D}$ increase with $P$, over a decade, resulting in two clearly defined but distinct slopes which will be analyzed in-depth in the Sec. 4. In contrast, the effective Coulomb friction coefficient $\mu_{D}$ shown in Fig. 4b appears weakly dependent on the confinement pressure and the particle properties, with an average value $\langle\mu_{D}\rangle\simeq 0.28$.
Finally, we look at the energetical performances of the granular packing, as defined in Eqs. 6-8; again, we normalize the energies per unit volume by $k_{n}$ (i.e. by the particle’s Young’s Modulus) for dimensional reasons. Fig. 4c shows that both the dissipated and the maximum stored energy per cycle increase as functions of the confinement pressure, consistently with the elastic features of the effective model, $G_{E}$ and $G_{D}$. Also in agreement with the effective Coulomb friction coefficient $\mu_{D}$, their ratio (i.e. the loss factor), shown in Fig. 4d, appears weakly dependent on the pressure and the particle properties, situated close to the value $\eta=0.40$ already observed in Fig. 3d.
Figure 4: (a): Dimensionless fitted parameters $G_{D}/k_{n}$ (open symbols) and $G_{E}/k_{n}$ (solid symbols) for various volume fractions $\phi$ and elasticities $E_{p}$ as functions of the dimensionless pressure $P/k_{n}$ at $\gamma_{m}=0.15$. (b): Dahl friction coefficient versus dimensionless pressure for the same data set. (c): Dissipated energy per cycle (open symbols) and maximum stored energy (solid symbols) computed from the hysteresis response using Eqs. 6 and 7 respectively, as functions of the dimensionless pressure. (d): Loss factor computed from 8, as a function of the dimensionless pressure. Legend symbols are common for (a-d).